### v1

# Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities

Abstract
Social scientists use surveys to probe the opinions and beliefs of populations, but these methods are slow, costly, and prone to biases. Recent advances in large language models (LLMs) enable creating computational representations or “digital twins” of populations that generate human-like responses mimicking the population’s language, styles, and attitudes. We introduce Community-Cross-Instruct, an unsupervised framework for aligning LLMs to online communities to elicit their beliefs. Given a corpus of a community’s online discussions, Community-Cross-Instruct automatically generates instruction-output pairs by an advanced LLM to (1) finetune an foundational LLM to faithfully represent that community, and (2) evaluate the alignment of the finetuned model to the community. We demonstrate the method’s utility in accurately representing political and fitness communities on Reddit. Unlike prior methods requiring human-authored instructions, Community-Cross-Instruct generates instructions in a fully unsupervised manner, enhancing scalability and generalization across domains. This work enables cost-effective and automated surveying of diverse online communities.
Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities
Zihao He, Rebecca Dorn, Siyi Guo, Minh Duc Chu, Kristina Lerman USC Information Sciences Institute {zihaoh, rdorn, siyiguo, mhchu}@usc.edu, lerman@isi.edu
1 Introduction
Social scientists use surveys and focus groups to learn the opinions, needs and concerns of diverse populations. However, designing surveys and recruiting participants is a slow and costly process, limiting the utility of these instruments for probing public opinion. Surveys are prone to biases, such as the social desirability bias Gordon (1987), where respondents may alter their responses to sensitive questions to appear more socially acceptable Bergen and Labonté (2020), non-response bias Hill et al. (1997), where participants fail to answer questions, and self-selection bias due to the choices participants make to participate in the survey Heckman (1990). In addition, social stigmas may taint responses Goel and Salganik (2010), especially for hard-to-reach and marginalized groups.
Figure 1: Illustration of Community-Cross-Instruct to align an LLM to a community. (1) Open-ended instructions and multi-choice survey questions are generated by an advanced LLM from the community data. (2) A foundational LLM is aligned to community through instruction-tuning on the open-ended instructions. (3) The alignment of the finetuned LLM to the community is measured on the survey questions.
Recent breakthroughs in generative AI and especially large language models (LLMs) enable new capabilities for creating computational representations of human populations — their digital twins El Saddik (2018) — by ingesting vast textual data they create, for example, in online discussion forums. These LLM-based models generate human-like responses that mimic the language, communication styles, and attitudes of populations they are aligned to, allowing us to probe their worldviews, biases, and sentiments in a cost-effective and automated manner. Previous works have leveraged such LLM-based representations to mine opinions and learn political attitudes of online communities Jiang et al. (2020); He et al. (2024b). However, they finetune GPT-2 Radford et al. (2019) directly on textual data from communities, where the finetuned model is not able to answer questions that may be in different format, e.g., multiple choice, as it is not an instruction-tuned model.
To address these challenges, we introduce Community-Cross-Instruct: a fully unsupervised framework for aligning LLMs to online communities through instruction-tuning. The finetuned LLMs serve as digital twins of communities, which can be automatically surveyed to elicit their views (Figure 1).
Specifically, given a corpus of comments and submissions in different forums on Reddit, Community-Cross-Instruct uses an advanced LLM (Claude-3) to automatically curate (1) CommInst: a set of open-ended instructions (Fig. 2(a)) for community-specific instruction tuning, and (2) CommSurvey, a set of multi-choice survey questions (Fig. 2(b)). Each instruction in CommInst and each survey question in CommSurvey are paired with answers from different communities. We finetune foundational LLMs (GPT-3.5 or Llama-3) on CommInst, in order to align them to different communities, and evaluate the finetuned LLMs on CommSurvey to measure alignment. See Figure 3 for a high-level overview of the framework. Existing methods for self-supervised instruction generation rely on a seed set of instructions curated by domain experts Wang et al. (2022); Chen et al. (2024b), which is not generalizable. We overcome this bottleneck by incorporating the readily available community data into the instruction generation pipeline. Our Community-Cross-Instruct framework distills knowledge from communities using an advanced LLM to generate instructions for aligning foundational LLMs.
Our key contributions can be summarized as
  • We introduce Community-Cross-Instruct, a novel unsupervised framework for aligning foundational LLMs to online communities, by finetuning them on the automatically curated set of open-ended community-specific instruction-answer pairs (CommInst). The models’ alignment to communities is measured using another set of automatically generated multiple choice questions and answers (CommSurvey).
  • Using data from Reddit forums, we show that our method improves the fidelity of community representation (alignment) in two domains—politics and fitness—yielding significant alignment improvement over standard persona adaptation methods.
Our work highlights the potential of generative AI to help researchers gain insights from online communities. By leveraging LLMs to create digital twins of these communities, researchers can more accurately and efficiently understand the nuances of public opinion, attitudes, and behaviors. Our framework not only enhances the fidelity of community representation but also paves the way for more effective approaches to studying social phenomena in the digital age.
Clarification on contributions.
We aim to align LLMs to online communities through instruction-tuning, so that the aligned LLMs can be flexibly surveyed to probe the communities’ views on different topics. To prepare data for instruction tuning, we propose Community-Cross-Instruct, and the automatic creation of the open-ended instruction-answer pairs (CommInst) is the core contribution. However, in order to demonstrate that the finetuned LLMs are indeed aligned to communities and can be treated as their digital twins, we propose to automatically generate survey questions and answers (CommSurvey) to efficiently evaluate LLM alignment. This is a secondary contribution of this paper, and we do not claim that CommSurvey is the best way to measure alignment. Therefore, our main goal is to propose a method to prepare LLMs as a proxy for surveying different online communities, rather than automatically creating such surveys.
On the use of the term alignment.
Inspired by Santurkar et al. (2023), throughout this paper, we use alignment to refer only to the alignment of views and opinions of LLMs and humans.
2 Problem Definition
A topical domain (e.g., politics or fitness) includes $n$ communities {$C_{1}$, $C_{2}$, …., $C_{n}$} each with different views and beliefs. Members of each community $C_{i}$ collectively author text corpus $D_{i}$ (e.g., discussions on Reddit forums) expressing views and exhibiting behaviors. Our goal is to align an LLM $f$ to each community $C_{i}$ using its texts $D_{i}$, such that the aligned LLM $f^{\prime}_{i}$ learns the complex mindset of the community and responds to inputs in the community’s voice. By administering surveys to the aligned LLMs $\{f^{\prime}_{1},f^{\prime}_{2},...,f^{\prime}_{n}\}$, we obtain responses from different communities, thereby capturing their ideological differences.
To align an LLM $f$ to a community $C$, we finetune it on a set of demonstrations (instruction-answer pairs) Wang et al. (2023); Ouyang et al. (2022); Chen et al. (2024b) $I=\{(X_{j},Y_{j})\}_{j=1}^{m}$, where the instructions are open-ended questions probing the community’s views on different topics, and the corresponding answers are aligned with each community’s ideology. Figure 2(a) shows an example demonstration in the politics domain. We propose Community-Cross-Instruct (Fig. 3), a framework to automatically generate community-specific demonstrations $I$ with an advanced LLM $\hat{f}$ (Claude-3) based on the community’s text corpus $D$ and use these demonstrations to instruction-tune a foundational LLM $f$ (GPT-3.5 or Llama-3) through a process we call “Cross-Instruct”.
Figure 2: Example of (a) an instruction from CommInst-Pol and (b) a survey question from CommSurvey-Pol. The instruction/survey question is paired with answers from different communities.
Figure 3: Overview of Community-Cross-Instruct, with an illustrative example of the politics domain. (1) Data collection for each community within the desired domain. (2) BERTopic clusters the data and identifies prominent topics. (3) For each topic, the advanced LLM is prompted with (i) on-topic examples from each community and (ii) instruction-answer task definition and explanation. (4) Creation of (a) open-ended instruction-answer pairs and (b) multi-choice survey questions and answers, for fine-tuning and evaluating a sub-population representation foundational LLM respectively.
3 CommInst: A Collection of Community-Specific Instructions
To align LLMs to different communities within each domain, we curate a dataset CommInst for community-specific instruction tuning. CommInst-Pol and CommInst-Fit are the two subsets for the politics and fitness domains with 6,687 and 4,395 unique instructions respectively. Each instruction is an opinion-eliciting question about a topic relevant to the domain (Figure 2(a)), which appears in multiple demonstrations. Each demonstration’s instruction is paired with an open-ended answer from a different community.
3.1 Online Community Corpora Collection
Reddit is a vibrant social media platform hosting discussion forums (subreddits) on a wide range of topics Hofmann et al. (2022); Chen et al. (2024a). In this paper we focus on two domains: politics and fitness, which contain distinctive subreddits, or communities, with complex social dynamics. Political subreddits are valuable expressions of diverse public opinions and viewpoints. In the age of rising polarization, LLMs can help researchers track the complex evolving ideological landscape and effectively elicit public opinions. Meanwhile, the fitness subreddits feature a wealth of sensitive, health-related conversations, and the latest diet and fitness fads, often discussed using inscrutable insider jargon. We can track the direction of these discussions, identifying emerging trends, risks, and potential health misinformation. This insight allows public health officials to promptly address harmful health advice and intervene where necessary. Additionally, it provides an early warning system to detect the spread of dangerous health practices, ensuring that corrective measures can be implemented swiftly to protect community health.
We identify a set of representative subreddits for each domain based on personal knowledge and by querying ChatGPT. We manually aggregate and filter results, obtaining the following political online communities: r/Anarchism, r/Anarcho_Capitalism, r/Conservative, r/Libertarian, r/Republican, and r/democrats; For fitness, we investigate: r/Fitness, r/loseit, r/keto, r/xxfitness, and r/GettingShredded. More details of these communities are presented in Appendix A.1.
We collect comments and submissions from January 2019 to December 2022 and remove those that were deleted by the author or the moderator. The statistics of the collected comments and submissions are shown in Appendix A.1. To balance the number of comments and submissions across different communities, for each community, we keep a maximum of 140K comments and 30K submissions. Therefore, there are at most 170K texts, which makes up the text corpus $D$, including comments and submissions, for each community.
3.2 Topic Modeling
For a domain, denote the combined corpus from all $n$ communities by $\mathbb{D}=D_{1}\cup D_{2}\cup....\cup D_{n}$. We use BERTopic Grootendorst (2022) on $\mathbb{D}$ to identify topics $g$. More details about BERTopic can be found in Appendix B.1. After topic modeling, each text $d$ is assigned a topic $g(d)$. The number of topics detected in politics and fitness are 425 and 628 respectively. The top 10 most frequent topics in politics and fitness are shown in Appendix B.1 respectively. Politics topics include abortion, climate, guns, and racism. Fitness topics include body measures, keto diet, and weight lifting.
For each community $C_{i}$, we split the text corpus $D_{i}$ into smaller chunks $D_{i}=\{\hat{D}^{i}_{1},\hat{D}^{i}_{2},....\}$, where each chunk $\hat{D}$ consists of texts belonging to a single topic $g(\hat{D})$.
3.3 Open-ended Instruction-Answer Pair Generation
For a specific domain, starting with an empty demonstration pool, we randomly sample without replacement a text chunk $D_{i}$ for each topic $t_{k}\in T$ from each community. We include the sampled text chunks on topic $t_{k}$ from the $n$ communities into a prompt and then use the prompt to query the advanced LLM to generate open-ended instructions and answers on the topic. The prompting template is shown in Appendix B.2. We query claude-3-opus-20240229 using the Anthropic API. In the prompt, we first specify the topic by the topic keywords, and then the comments in the sampled chunks from different communities. Next we ask the model to generate instructions that can be answered based on the texts, and that the instructions should elicit different responses from the $n$ communities. Each generated instruction is paired with different answers from different communities (Figure 2(a)). It is then added to the pool as $n$ demonstrations. Note that some topics may be discussed more in certain communities than others, so some instructions may be paired with answers from less than $n$ communities. We iteratively repeat this process until there are no text chunks left for topic $t_{k}$ from any community, until we shift to the next topic $t_{k+1}$. After the iterative generation process, the instruction-answer pool will contain demonstrations covering various topics, which will use to finetune foundational LLMs. Please refer to Appendix A.2 for more details about CommInst.
For fair evaluating the finetuned LLMs, we split CommInst-Pol in two ways:
  • Random split. CommInst-Pol is randomly split into training and eval sets: CommInst-Pol-train and CommInst-Pol-eval, in a ratio of 85:15.
  • Topical split. We randomly select 10 topics in politics and collect demonstrations related to the topics into CommInst-Pol-eval-topic. The remaining demonstrations are collected as CommInst-Pol-train-topic. In this way, CommInst-Pol-train-topic and CommInst-Pol-eval-topic has no topic overlap.
We repeat the same process for CommInst-Fit.
4 CommSurvey: A Collection of Community-Specific Survey Questions
To measure the alignment of the finetuned foundational LLM to the community, we administer a survey of multi-choice questions to the model and evaluate the agreement between the model’s responses and the community members’ responses. However, manually designing surveys and collecting human responses from online communities is a non-trivial process, which is costly and time-consuming. Instead, we curate another dataset called CommSurvey for evaluating the alignment of the finetuned LLM’s views to the community using survey questions. CommSurvey consists of automatically generated multi-choice survey questions and answers, where each question is paired with answers from different communities (Figure 2(b)). CommSurvey-Pol and CommSurvey-Fit are the two subsets for politics and fitness respectively, consisting of 4,457 and 2,930 multi-choice questions respectively, covering a wide range of topics.
politics r/Anarchism
r/Anarcho_
Capitalism
r/Conservative r/Libertarian r/Republican r/democrats
0.75 0.60 0.75 0.45 0.75 0.65
fitness r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
0.75 0.70 0.60 0.50 0.65
Table 1: Agreement between human annotators’ survey responses and the advanced LLM (Claude-3) generated survey answers in different domains and subreddits, measured by accuracy.
4.1 Multi-choice Survey Generation.
We initialize an empty multi-choice survey pool. Then, following the open-ended instruction-answer pairs generation in §$\S$3.3, we iteratively generate multi-choice questions and answers (Figure 2(b)) using the same advanced LLM (Claude-3) and add them to the pool. We also explored other advanced LLMs, such as GPT-4 or Gemini, but none of them was able to generate high-quality questions or answers as Claude-3. The prompting template is shown in Figure 8. We assume that the answers generated by the advanced LLM are the “semi-ground truths”, and that they faithfully represent the views of the corresponding communities.
Note that there is a correspondence between instructions in CommInst and survey questions CommSurvey, as they are generated using nearly identical prompts. Following the correspondence, we split CommSurvey-Pol and CommSurvey-Fit following the same two ways for CommInst. Due to the correspondence, there is no information leakage between CommInst-Pol-train and CommSurvey-Pol-eval, or between CommInst-Pol-train-topic and CommSurvey-Pol-eval-topic.
4.2 Validating Generated Survey Answers
We verify the faithfulness of the “semi-ground truths” for the communities via human annotation. For each domain, we randomly sample 20 questions from the survey (see Fig. 9 and Fig. 10 in the Appendix) and ask two expert annotators to answer them from the perspective of the subreddit. We also tried directly posting the surveys to the subreddits and having community members answer them. However, our posts were immediately removed by moderators and our accounts were banned from certain subreddits. This further indicates the difficulty of directly surveying online communities.
[…]
5 Experiments
We demonstrate the effectiveness of the pipeline on the Reddit politics and fitness domains. We align LLMs to Reddit communities by finetuning them on relevant demonstrations from CommInst. After finetuning, we administer surveys from CommSurvey and evaluate the finetuned LLMs by comparing their responses to the “semi-ground truths”, in order to measure alignment.
5.1 Experimental Setup
We experiment with two foundational LLMs – GPT-3.5 and Llama-3.
Finetuning.
We finetune the foundational LLMs on CommInst-Pol-train and CommInst-Fit-train. We use the OpenAI finetuning API to finetune gpt-3.5-turbo-0125 for 1 epoch, as we find that the loss converges within 1 epoch. The batch size is automatically determined by the API, and the finetuning takes ~40mins. For Llama-3, the implementation is based on LlamaFactory Zheng et al. (2024), and we finetune Llama-3-8B-Instruct with batch size 16 for 2 epochs on 4 H100 GPUs. The finetuning takes ~10mins.
Measuring Alignment.
We administer survey questions in CommSurvey-Pol-veal and CommSurvey-Fit-veal to the finetuned foundational LLMs. In the prompt, we provide the question and the four options verbatim, and then add the instruction “Select only one answer by stating either A, B, C, or D. Do not provide any additional explanation or rationale for your choice.”, for the sake of easier matching of the model’s response to the option. For GPT-3.5, the temperature is set to 0.9 and the model provides 5 generations for each question. For Llama-3, temperature is set to 0.9, $top\_p$ is set to 0.9, and the model provides 10 generations for each question. The majority vote from the different generations for each question is taken. We compare the model responses to the “semi-ground truths” and report the accuracy as a measure of alignment.
5.2 Baselines
Vanilla LLMs.
We administer survey questions to the vanilla unfinetuned GPT-3.5 and Llama-3 and evaluate their answers.
Vanilla LLMs+Steering.
When prompting the vanilla language models to answer survey questions, we steer them to mimic the community, by specifying in the prompt that “Select only one answer that best aligns with the opinions of members from subreddit r/[subreddit]”. Steering the vanilla LLMs can nudge them to respond for the community. However, it is important to note that steering only applies to predefined communities that are developed via a predefined tag, such as subreddits, where we can easily specify community names in the prompt to steer the LLMs. Although in this paper we focus on communities associated with online forums, Community-Cross-Instruct generalizes to other online communities, including organically formed communities in the retweet network Chu et al. (2024) or the news co-sharing network Mosleh et al. (2021); He et al. (2024b), as long as their texts are readily available. It is non-trivial to summarize the organically-formed communities concisely in language, but we can still align LLMs to them via Community-Cross-Instruct.
5.3 Main Results
We compare survey answers generated by the finetuned LLMs to the “semi-ground truth” and report the accuracy in Figure 4 as a measure of alignment with the corresponding community. In most cases, steering LLMs towards specific communities creates better alignment by relying on the models’ pre-existing knowledge about the communities. However, such pre-knowledge may be incomplete, inaccurate, or outdated. Therefore, by finetuning LLMs on instructions distilled from the text corpora of communities, Community-Cross-Instruct outperforms steering and achieves the best results, demonstrating its effectiveness in aligning LLMs to different communities.
(a) politics
(b) fitness.
Figure 4: Alignment of finetuned LLMs to the corresponding communities measured by accuracy, by comparing the LLMs’ survey responses to the “semi-ground truths”.
The base models show worse alignment with r/Conservative and r/Republican than other Reddit communities. This observation conforms to findings by previous works Santurkar et al. (2023); He et al. (2024a); Chen et al. (2024b) that LLMs tend to display a liberal bias and poor alignment with conservatives. After explicitly steering them towards specific communities, their alignment with the two subreddits significantly increases, but it is still lower than that with other communities. Finetuning the LLMs with Community-Cross-Instruct further improves the alignment, although it is still lower than for other communities. This observation highlights the risks of deploying LLMs in political scenarios due to the entrenched political biases.
For the r/Anarchism subreddit, GPT-3.5-CrossInst achieves similar alignment as the steering baseline. We suspect that this is because views about anarchism are captured by the pre-existing knowledge of GPT-3.5. Therefore, infusing such knowledge again by Community-Cross-Instruct barely leads to any gain.
For the r/Fitness and r/xxfitness communities, steering both LLMs worsens the alignment compared to the vanilla model. r/Fitness is a general fitness subreddit and r/xxfitness is a fitness community for women, both of which may have a broader and more diverse range of opinions compared to other subreddits. Steering the models to align with a singular “community opinion” might oversimplify the nuanced perspectives within these fitness communities, leading to decreased alignment accuracy. However, Community-Cross-Instruct effectively distills knowledge from the community corpora during finetuning. On r/GettingShredded, Community-Cross-Instruct is outperformed by the steering baseline for GPT-3.5, but not Llama-3. r/GettingShredded has the least amount of comments and posts, producing much fewer demonstrations for finetuning (Appendix 3). The insufficient data might be less effective for finetuning a larger LLM like GPT-3.5111We assume that GPT-3.5 has more than 8B parameters., leading to inferior performance.
5.4 Out-of-Topic Generalizability
Our final goal is to create LLMs aligned to different communities, which can be used to answer any survey question from the perspective of the community. In the real world, the surveys may contain questions covering topics that do not appear in the finetuning data CommInst. In this study, we finetune Llama-3 on CommInst-Pol-train-topic, and evaluate it on CommSurvey-Pol-eval-topic, to make sure that there is no overlap in topic coverage. The results in Fig. 5 show that Community-Cross-Inst consistently outperforms both baselines, demonstrating its superiority in aligning with views of different communities, and generalize the learned views to unseen topics.
Figure 5: Cross-topic alignment of finetuned LLMs to the corresponding communities. The LLMs are finetuned on demonstrations and evaluated on survey questions on different topics.
6 Related Works
Self-Improved Instruction Tuning. A survey of instruction-tuning approaches Zhang et al. (2023) outlines several methods for models to autonomously self-improve their instruction set. These include generating instructions for pre-existing texts Li et al. (2023), eliciting interaction between different model iterations Chen et al. (2024c), and bootstrapping from an existing instruction set to generate new ones Wang et al. (2022). Additionally, a human-in-the-loop framework builds upon self-generalization by iterating between human and machine-generated instructions Guo et al. (2024). Despite these advances, these strategies require manual input. In this work, we build on self-instruct, a method to enhance instruction tuning by eliciting synthetic model instruction generations. We alter the original framework by removing the required set of manually written seed instructions Wang et al. (2022). This work significantly contributes to the field by offering a scalable and fully autonomous solution to instruction tuning, paving the way for more adaptive and intelligent models.
Community Alignment. Aligning with the dialects and behaviors of online communities is crucial for accurately modeling subpopulations Dorn et al. (2023, 2024). Subpopulation representative models (SRMs) Simmons and Hare (2023) can be used to emulate some characteristics of a particular subpopulation, particularly as LLMs can provide fine-grained, demographically-correlated outputs Argyle et al. (2023a). Using self-instruct with questions from the American Trans Panel as the seed instruction set, LLMs can be ideologically manipulated to reflect particular political opinions Chen et al. (2024b). To learn about partisan communities, Jiang et al. (2022) repeatedly prompt GPT 2 with tweets authored by prominent community members to generate synthetic responses to election survey questions. Argyle et al. (2023b) find that exposing GPT 3 to thousands of socio-demographic backstories leads the model to obtain a complex understandings of sociological concepts. ‘Media diet models’ train BERT using MLM on news articles, evaluating the alignment of model predictions with a survey of community members who follow the media Chu et al. (2023). Feldman et al. (2022) expose GPT 3 to Yelp reviews to expand insights past sentiment information. Prompting language models with (survey question, individual ID, survey year) triplets from the Americans in the General Social Surveys has helped social scientists impute missing survey data and predict opinions Kim and Lee (2023). These techniques rely on manually created datasets at least as a seed dataset to align the models with the desired communities, limiting SRBMs to high-resource environments.
Surveys. Social scientists use surveys to systematically collect data from populations to characterize their beliefs, attitudes, opinions, and behaviors. While designing the survey, researchers consider the order, wording, and format of questions (open-ended vs multiple-choice) to increase clarity and ensure the validity and reliability of the survey data. Survey design also aims to minimize biases, such as the non-response bias Hill et al. (1997), fatigue due to overly long surveys Choi and Pak (2005), social desirability bias Gordon (1987) when asking survey participants questions about sensitive topics (due to the desire to conform to perceived social expectations), or those involving social stigmas Goel and Salganik (2010). These issues, as well as the variety of existing biases, call for new approaches to survey design.
7 Conclusion
In this paper, we introduced and explored Community-Cross-Instruct, a fully automated framework to represent online communities and elicit their views. By automating the process of surveying diverse online communities, Community-Cross-Instruct offers a cost-effective and efficient alternative to traditional survey methods used by social scientists. Through finetuning LLMs based on community-specific instructions and answers, Community-Cross-Instruct enables the generation of accurate responses that align with the beliefs and perspectives of different online communities. This innovative framework has demonstrated its effectiveness in representing various communities, including political and fitness forums on Reddit. Community-Cross-Instruct opens up new possibilities for researchers to gain insights into the diverse views and opinions present in online communities, paving the way for deeper understanding and analysis of digital societies. Moving forward, we will apply the framework to model organic online communities, such as those in the retweet network.
Limitations
Identical LLM for Generating both Datasets.
We use Claude-3 for generating both CommInst for finetuning LLMs and CommSurvey for evaluating LLM alignment, as we couldn’t find a different advanced LLM that was able to generate high-quality data as Claude-3. This reliance on a single LLM introduces a limitation, as it may lead to a bias where the evaluation dataset (CommSurvey) is inherently aligned with the model used for instruction generation (CommInst), potentially overestimating the alignment accuracy of the finetuned models. To address this limitation, we plan to incorporate a diverse set of advanced LLMs as they become available in future iterations of our framework. Additionally, we will include human-in-the-loop validation to ensure the generated datasets maintain high quality and representativeness, mitigating any biases introduced by the single LLM dependency. Furthermore, we will explore cross-validation techniques and third-party evaluations to benchmark the performance of our finetuned models, ensuring robustness and generalizability of our results beyond the influence of a single LLM.
Group Approximation.
To approximate group-level behavior, community members represented in the minority are inherently excluded. Further, measuring group alignment using a single-answer multiple choice questionnaire does not account for bimodal distribution. We hope to build on this work by experimenting with survey designs that account for more diversity in communities.
Hallucination Potential.
Adapting models to communities poses a risk of language models hallucinating or providing misinformation in their community representation. We hope to build upon this work with in-depth experiments on model hallucination and misrepresentation in the subgroup representation task.
Prompt Perturbations.
LLM responses are sensitive to slight changes in prompts Salinas and Morstatter (2024). In this work, we work primarily with one prompt for instruction generation. We would be interested to see how model’s generated instructions shift with different prompting schemas.
Ethics Statement
Users might not be informed of how their data (reflected in their posts) are used to produce digital twins that mirror their voices and the purpose of the survey constructed based on their data. Furthermore, automatically surveying communities can reveal unconsented insights of certain individuals or groups of people online. To address these ethical concerns, we will implement several measures. Firstly, we will ensure that all data used for creating digital twins is anonymized, stripping any personally identifiable information to protect user privacy. Secondly, we will seek to aggregate data in a manner that focuses on community-level insights rather than individual-level analysis, thereby reducing the risk of unconsented personal exposure. Thirdly, we will obtain explicit consent from users where possible, clearly communicating the purpose and scope of our research. Lastly, we will adhere to ethical guidelines and institutional review board (IRB) requirements to ensure that our methods respect the privacy and rights of the individuals and communities involved.
Appendix A Data Details
A.1 Subreddits
We briefly introduce each subreddit in Table 2. Statistics of data from different subreddits are shown in Table 3.
Subreddit Description
r/Anarchism
r/Anarchism is a forum with 268K members for discussing anarchism, assuming
participants are anarchists, allies, or learners.
r/Anarcho_Capitalism
r/Anarcho_Capitalism is a subreddit with 197K members dedicated to discussing
free-market capitalist anarchism and related topics, advocating for a society where
voluntary interactions enhance liberty and opportunity for all.
r/Conservative
r/Conservative, with 1.1M members, offers the largest space on Reddit for fiscal
and social conservatives to explore and discuss political and cultural issues from
a distinctly conservative perspective.
r/Libertarian
r/Libertarian is a forum with 506K members focused on libertarianism and the
principles of a free society, explicitly rejecting anti-libertarian ideologies
like communism.
r/Republican
r/Republican is a partisan subreddit with 191K members specifically designed
as a space for Republicans to discuss issues among themselves.
r/democrats
r/democrats is a subreddit with 438K members dedicated to supporting the
Democratic Party, offering daily updates, policy analysis, and opportunities to
help elect Democrats at all levels.
r/Fitness
r/Fitness is the largest community dedicated to achieving physical fitness
goals with 12M members
r/loseit
r/loseit is a supportive community with 4M members for individuals of all
sizes to discuss healthy and sustainable weight loss methods
r/keto
r/keto is a community of 3.6M members for sharing experiences and advice
about the low-carb Ketogenic diet, which supports a range of health
conditions from diabetes to epilepsy
r/xxfitness
r/xxfitness is a community of 2.2M members for female and gender
non-binary/gender non-conforming Redditors to discuss fitness
r/GettingShredded
r/GettingShredded is a community of 162K members for anyone aiming to lose
body fat, improve their physique, and boost confidence through healthy fat loss
and strength training discussions.
Table 2: Reddit forums used in this study. Descriptions of (top) politics and (bottom) fitness subreddits.
A.2 Statistics of CommInst
Table 3 describes the number of demonstrations in each subreddit. The variance is due to their different topic coverage. Figure 6 presents the distribution of instruction lengths and answer lengths in CommInst-Pol and CommInst-Fit.
Politics
r/Anarch. r/Anarcho. r/Conserv. r/Democr. r/Libert. r/Republic.
Comments 182,192 589,354 2,391,745 142,274 1,950,089 140,572
Submissions 31,567 41,264 356,842 90,001 40,967 48,808
Demonstrations 2,430 3,027 2,943 3,159 2,928 2,574
Fitness
r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
Comments 907,839 967,645 603,920 420,082 78,053
Submissions 2,430 3,027 2,943 3,159 2,928
Demonstrations 3,438 2,793 1,716 2,883 1,215
Table 3: Dataset size measured by number of comments and submissions for each community in political and fitness communities, and the number of learned open-ended instruction-answer pairs (demonstrations).
(a) Instructions in CommInst-Pol
(b) Instructions in CommInst-Fit
(c) Answers in CommInst-Pol
(d) Instructions in CommInst-Fit
Figure 6: Distribution of instruction and answer lengths in CommInst-Pol and CommInst-Fit.
Appendix B Implementation Details
B.1 Topic Modeling with BERTopic
Following the pipeline of BERTopic Grootendorst (2022), we first obtain the text embeddings using all-mpnet-base-v2 Song et al. (2020). The embedding dimensions are reduced with UMAP, with n_neighbors 15 and n_components 5. Then the embeddings are clustered using HDBSCAN, with min_cluster_size 50. After fitting the model, we remove texts that are not assigned any topics. Table 4 and Table 5 present the top-10 most frequent topics in politics and fitness.
idx Topic Keywords Frequency
1 covid, vaccinated, masks, vaccines, healthcare, deaths, flu, pandemic, coronavirus, cdc 48,787
2 abortion, fetus, abortions, roe, body, choice, unborn, rights, parenthood, cells 11,256
3 climate, change, solar, emissions, pipeline, fuels, coal, power, supply, grid 9,035
4 guns, firearms, shootings, firearm, amendment, laws, armed, rifles, militia, nra 8,833
5 police, officer, officers, gun, force, law, crime, video, training, situation 7,934
6 kamala, woman, pelosi, vp, hillary, president, herself, biden, party, with 7,879
7 racist, racism, whites, racists, racial, minorities, democrats, privilege, whiteness, discrimination 7,785
8 libertarian, libertarians, libertarianism, party, government, left, ideology, vote, political, conservative 6,666
9 anarchism, anarchy, anarchist, anarchists, anarcho, hierarchies, communism, ideology, government, system 6,311
10 biden, bernie, obama, candidate, voters, debate, election, democrats, hillary, sanders 5,768
Table 4: Top-10 most frequent topics in politics. Each topic is represented by the topic-10 keywords.
idx Topic Keywords Frequency
1 lbs, bmi, down, cw, gained, progress, healthy, tips, weeks, started 18,296
2 keto, ketosis, carbs, diet, 20g, protein, macros, started, calories, ketogenic 13,078
3 leggings, clothes, wear, pants, shorts, bras, shirts, clothing, dress, lululemon 7,063
4 motivation, journey, progress, yourself, change, motivated, success, start, proud, back 5,142
5 water, scale, weighing, retention, trend, normal, plateau, fluctuates, gain, weekly 4,880
6 sleep, morning, hours, work, gym, motivation, schedule, waking, exercise, routine 4,486
7 she, people, overweight, weight, obese, comment, healthy, skinny, talk, shaming 4,192
8 bf, percentage, fat, bodyfat, body, lean, bulk, estimates, lbs, pic 3,930
9 reps, sets, program, max, amrap, lifts, ranges, bench, 5x10, hypertrophy 3,788
10 shoulder, elbow, injury, arm, knees, physio, exercises, bicep, curls, heal 3,775
Table 5: Top-10 most frequent topics in fitness. Each topic is represented by the topic-10 keywords.
B.2 Prompting Template for CommInst
The prompting template for generating the open-ended instruction-answer pairs in CommInst is shown in Figure 7.
Figure 7: Template prompt to generate CommInst.
B.3 Prompting Template for Generating CommSurvey
The prompting template for generating the multiple choice instruction-answer pairs in CommSurvey is shown in Figure 8.
Figure 8: Template prompt to generate CommSurvey.
Appendix C Survey Sample
Samples of the surveys provided to human annotators for generating “semi-ground truth data” are provided in Figures 9 and 10 (political and fitness domains, respectively).
Figure 9: Sample of the multiple choice survey generated for human evaluation in the politics domain.
Figure 10: Sample of the multiple choice survey generated for human evaluation in the fitness domain.

### v2

@@
 Social scientists use surveys to probe the opinions and beliefs of populations, but these methods are slow, costly, and prone to biases.
-Recent advances in large language models (LLMs) enable creating computational representations or “digital twins” of populations that generate human-like responses mimicking the population’s language, styles, and attitudes.
+Recent advances in large language models (LLMs) enable the creation of computational representations or “digital twins” of populations that generate human-like responses mimicking the population’s language, styles, and attitudes.
 We introduce Community-Cross-Instruct, an unsupervised framework for aligning LLMs to online communities to elicit their beliefs.
-Given a corpus of a community’s online discussions, Community-Cross-Instruct automatically generates instruction-output pairs by an advanced LLM to (1) finetune an foundational LLM to faithfully represent that community, and (2) evaluate the alignment of the finetuned model to the community.
-We demonstrate the method’s utility in accurately representing political and fitness communities on Reddit.
+Given a corpus of a community’s online discussions, Community-Cross-Instruct automatically generates instruction-output pairs by an advanced LLM to (1) finetune a foundational LLM to faithfully represent that community, and (2) evaluate the alignment of the finetuned model to the community.
+We demonstrate the method’s utility in accurately representing political and diet communities on Reddit.
 Unlike prior methods requiring human-authored instructions, Community-Cross-Instruct generates instructions in a fully unsupervised manner, enhancing scalability and generalization across domains.
-This work enables cost-effective and automated surveying of diverse online communities.
+This work enables cost-effective and automated surveying of diverse online communities [^1] Code and data are available at https://github.com/zihaohe123/community-cross-instruct.
 Community-Cross-Instruct: Unsupervised Instruction Generation for Aligning Large Language Models to Online Communities
-Zihao He, Rebecca Dorn, Siyi Guo, Minh Duc Chu, Kristina Lerman USC Information Sciences Institute {zihaoh, rdorn, siyiguo, mhchu}@usc.edu, lerman@isi.edu
+Zihao He, Minh Duc Chu, Rebecca Dorn, Siyi Guo, Kristina Lerman USC Information Sciences Institute {zihaoh, mhchu, rdorn, siyiguo}@usc.edu, lerman@isi.edu
+**footnotetext: Equal Contribution.
 1 Introduction
-Social scientists use surveys and focus groups to learn the opinions, needs and concerns of diverse populations.
+Social scientists use surveys and focus groups to learn the opinions, needs, and concerns of diverse populations.
 However, designing surveys and recruiting participants is a slow and costly process, limiting the utility of these instruments for probing public opinion.
@@
 (1) Open-ended instructions and multi-choice survey questions are generated by an advanced LLM from the community data.
-(2) A foundational LLM is aligned to community through instruction-tuning on the open-ended instructions.
-(3) The alignment of the finetuned LLM to the community is measured on the survey questions.
+(2) A foundational LLM is aligned to the community through instruction-tuning on the open-ended instructions.
+(3) The alignment of the finetuned LLM to the community is measured using the generated survey questions.
 Recent breakthroughs in generative AI and especially large language models (LLMs) enable new capabilities for creating computational representations of human populations — their digital twins El Saddik (2018) — by ingesting vast textual data they create, for example, in online discussion forums.
@@
 (2024b).
-However, they finetune GPT-2 Radford et al.
-(2019) directly on textual data from communities, where the finetuned model is not able to answer questions that may be in different format, e.g., multiple choice, as it is not an instruction-tuned model.
-To address these challenges, we introduce Community-Cross-Instruct: a fully unsupervised framework for aligning LLMs to online communities through instruction-tuning.
-The finetuned LLMs serve as digital twins of communities, which can be automatically surveyed to elicit their views (Figure 1).
-Specifically, given a corpus of comments and submissions in different forums on Reddit, Community-Cross-Instruct uses an advanced LLM (Claude-3) to automatically curate (1) CommInst: a set of open-ended instructions (Fig.
-2(a)) for community-specific instruction tuning, and (2) CommSurvey, a set of multi-choice survey questions (Fig.
-2(b)).
-Each instruction in CommInst and each survey question in CommSurvey are paired with answers from different communities.
-We finetune foundational LLMs (GPT-3.5 or Llama-3) on CommInst, in order to align them to different communities, and evaluate the finetuned LLMs on CommSurvey to measure alignment.
-See Figure 3 for a high-level overview of the framework.
+However, these studies typically finetune models like GPT-2 Radford et al.
+(2019) directly on raw textual data from the communities, resulting in models that can perform text continuation but are limited in their ability to answer questions in structured formats, such as multiple choice.
+Moreover, raw community data often contains noise, irrelevant information, or off-topic discussions that can degrade model performance if not properly handled.
+To address these challenges, we aim to align LLMs to communities by finetuning the models with high-quality instructions where the community perspectives are embedded.
+However, curating high-quality instructional data is a non-trivial process.
 Existing methods for self-supervised instruction generation rely on a seed set of instructions curated by domain experts Wang et al.
 (2022); Chen et al.
-(2024b), which is not generalizable.
-We overcome this bottleneck by incorporating the readily available community data into the instruction generation pipeline.
-Our Community-Cross-Instruct framework distills knowledge from communities using an advanced LLM to generate instructions for aligning foundational LLMs.
+(2024b), which is limited by the generalizability to a new domain.
+In this paper, we introduce Community-Cross-Instruct: a fully unsupervised framework for aligning LLMs to online communities through instruction-tuning.
+It incorporates the readily available community data into the instruction generation pipeline, which no human supervision is needed.
+Community-Cross-Instruct uses advanced LLMs to generate instruction-response pairs that better capture community perspectives in more structured and useful formats.
+These instructions-responses pairs are used to finetune foundational LLMs.
+The finetuned LLMs serve as digital twins of communities, which can be automatically surveyed to elicit their views.
+We show a high-level overview of our framework in Figure 1.
+Specifically, given a corpus of comments and submissions in different forums on Reddit, Community-Cross-Instruct uses an advanced LLM (GPT-4o) to automatically curate two instructional datasets: (1) CommInst: a set of open-ended instructions for community-specific instruction tuning, and (2) CommSurvey, a set of multi-choice survey questions for community-specific survey completion.
+Each instruction in CommInst and each survey question in CommSurvey are paired with responses from different communities (Figure 2).
+We finetune foundational LLMs (GPT-3.5 or Llama-3.1) on CommInst, in order to align them to different communities and evaluate the finetuned LLMs on CommSurvey to measure alignment.
 Our key contributions can be summarized as
-  • We introduce Community-Cross-Instruct, a novel unsupervised framework for aligning foundational LLMs to online communities, by finetuning them on the automatically curated set of open-ended community-specific instruction-answer pairs (CommInst).
+  • We introduce Community-Cross-Instruct, a novel unsupervised framework for aligning foundational LLMs to online communities, by finetuning them on the automatically curated set of open-ended community-specific instruction-response pairs (CommInst).
 The models’ alignment to communities is measured using another set of automatically generated multiple choice questions and answers (CommSurvey).
-  • Using data from Reddit forums, we show that our method improves the fidelity of community representation (alignment) in two domains—politics and fitness—yielding significant alignment improvement over standard persona adaptation methods.
+  • Using data from Reddit forums, we show that our method improves the fidelity of community representation (alignment) in two domains—politics and diet—yielding significant alignment improvement over standard persona adaptation methods.
 Our work highlights the potential of generative AI to help researchers gain insights from online communities.
@@
 Our framework not only enhances the fidelity of community representation but also paves the way for more effective approaches to studying social phenomena in the digital age.
-Clarification on contributions.
-We aim to align LLMs to online communities through instruction-tuning, so that the aligned LLMs can be flexibly surveyed to probe the communities’ views on different topics.
-To prepare data for instruction tuning, we propose Community-Cross-Instruct, and the automatic creation of the open-ended instruction-answer pairs (CommInst) is the core contribution.
-However, in order to demonstrate that the finetuned LLMs are indeed aligned to communities and can be treated as their digital twins, we propose to automatically generate survey questions and answers (CommSurvey) to efficiently evaluate LLM alignment.
-This is a secondary contribution of this paper, and we do not claim that CommSurvey is the best way to measure alignment.
-Therefore, our main goal is to propose a method to prepare LLMs as a proxy for surveying different online communities, rather than automatically creating such surveys.
-On the use of the term alignment.
+Using the term alignment.
 Inspired by Santurkar et al. (2023), throughout this paper, we use alignment to refer only to the alignment of views and opinions of LLMs and humans.
 2 Problem Definition
-A topical domain (e.g., politics or fitness) includes $n$ communities {$C_{1}$, $C_{2}$, …., $C_{n}$} each with different views and beliefs.
+A topical domain (e.g., politics or diet) includes $n$ communities {$C_{1}$, $C_{2}$, …., $C_{n}$} each with different views and beliefs.
 Members of each community $C_{i}$ collectively author text corpus $D_{i}$ (e.g., discussions on Reddit forums) expressing views and exhibiting behaviors.
@@
 By administering surveys to the aligned LLMs $\{f^{\prime}_{1},f^{\prime}_{2},...,f^{\prime}_{n}\}$, we obtain responses from different communities, thereby capturing their ideological differences.
-To align an LLM $f$ to a community $C$, we finetune it on a set of demonstrations (instruction-answer pairs) Wang et al.
+To align an LLM $f$ to a community $C$, we finetune it on a set of demonstrations (instruction-response pairs) Wang et al.
 (2023); Ouyang et al.
 (2022); Chen et al.
-(2024b) $I=\{(X_{j},Y_{j})\}_{j=1}^{m}$, where the instructions are open-ended questions probing the community’s views on different topics, and the corresponding answers are aligned with each community’s ideology.
+(2024b) $I=\{(X_{j},Y_{j})\}$, where the instructions are open-ended questions probing the community’s views on different topics, and the corresponding responses are aligned with each community’s ideology.
 Figure 2(a) shows an example demonstration in the politics domain.
-We propose Community-Cross-Instruct (Fig.
-3), a framework to automatically generate community-specific demonstrations $I$ with an advanced LLM $\hat{f}$ (Claude-3) based on the community’s text corpus $D$ and use these demonstrations to instruction-tune a foundational LLM $f$ (GPT-3.5 or Llama-3) through a process we call “Cross-Instruct”.
-Figure 2: Example of (a) an instruction from CommInst-Pol and (b) a survey question from CommSurvey-Pol. The instruction/survey question is paired with answers from different communities.
+We propose Community-Cross-Instruct (Figure 3), a framework to automatically generate community-specific demonstrations $I$ with an advanced LLM $\hat{f}$ (GPT-4o) based on the community’s text corpus $D$ and use these demonstrations to instruction-tune a foundational LLM $f$ (GPT-3.5 or Llama-3.1) through a process we call “Cross-Instruct”.
+Figure 2: Example of (a) an instruction from CommInst and (b) a survey question from CommSurvey in the politics domain on the topic of marijuana.
+The open-ended instruction and survey question are paired with answers from different communities.
 Figure 3: Overview of Community-Cross-Instruct, with an illustrative example of the politics domain.
-(1) Data collection for each community within the desired domain.
+(1) Data is collected for each community within the desired domain.
 (2) BERTopic clusters the data and identifies prominent topics.
-(3) For each topic, the advanced LLM is prompted with (i) on-topic examples from each community and (ii) instruction-answer task definition and explanation.
-(4) Creation of (a) open-ended instruction-answer pairs and (b) multi-choice survey questions and answers, for fine-tuning and evaluating a sub-population representation foundational LLM respectively.
-3 CommInst: A Collection of Community-Specific Instructions
-To align LLMs to different communities within each domain, we curate a dataset CommInst for community-specific instruction tuning.
-CommInst-Pol and CommInst-Fit are the two subsets for the politics and fitness domains with 6,687 and 4,395 unique instructions respectively.
-Each instruction is an opinion-eliciting question about a topic relevant to the domain (Figure 2(a)), which appears in multiple demonstrations.
-Each demonstration’s instruction is paired with an open-ended answer from a different community.
-3.1 Online Community Corpora Collection
+A chunk is a set of documents from a community on the same topic.
+Chunki,j represents the chunk from community $i$ on topic $j$.
+(3) For each topic, the advanced LLM is prompted with (i) on-topic chunks from each community and (ii) task definition of the instructional data generation (see Appendix B.2), which leads the LLM to generate (a) open-ended instruction-response pairs and (b) multi-choice question-answer pairs.
+Ri,k represents the response of community $i$ to instruction $k$; Ai,k represents the answer of community $i$ to question $k$.
+(4) The open-ended instructions across all topics, along with the corresponding responses of community $i$, are added to CommInsti, which is used to finetune a foundational LLM, to align the LLM to the community.
+(5) The multi-choice questions across all topics, along with the corresponding answers from community $i$, are added to CommSurveyi, which is used to evaluate the finetuned LLM.
+3 Online Community Corpora Collection
 Reddit is a vibrant social media platform hosting discussion forums (subreddits) on a wide range of topics Hofmann et al.
@@
 (2024a).
-In this paper we focus on two domains: politics and fitness, which contain distinctive subreddits, or communities, with complex social dynamics.
+In this paper, we focus on two domains: politics and diet, which contain distinctive subreddits, or communities, with complex social dynamics.
 Political subreddits are valuable expressions of diverse public opinions and viewpoints.
 In the age of rising polarization, LLMs can help researchers track the complex evolving ideological landscape and effectively elicit public opinions.
-Meanwhile, the fitness subreddits feature a wealth of sensitive, health-related conversations, and the latest diet and fitness fads, often discussed using inscrutable insider jargon.
+Meanwhile, the diet subreddits feature a wealth of sensitive, health-related conversations, and the latest diet and diet fads, often discussed using inscrutable insider jargon.
 We can track the direction of these discussions, identifying emerging trends, risks, and potential health misinformation.
@@
 We identify a set of representative subreddits for each domain based on personal knowledge and by querying ChatGPT.
-We manually aggregate and filter results, obtaining the following political online communities: r/Anarchism, r/Anarcho_Capitalism, r/Conservative, r/Libertarian, r/Republican, and r/democrats; For fitness, we investigate: r/Fitness, r/loseit, r/keto, r/xxfitness, and r/GettingShredded.
+We manually aggregate and filter results, obtaining the following five political online communities: r/Liberal, r/NeutralPolitics, r/Anarcho_Capitalism, r/Conservative, and r/AskThe_Donald; For diet, we investigate three communities: r/keto, r/WeightLossAdvice, and r/EDAnonymous.
 More details of these communities are presented in Appendix A.1.
-We collect comments and submissions from January 2019 to December 2022 and remove those that were deleted by the author or the moderator.
+We collect comments and submissions from January 2019 to December 2022 and remove those that were deleted by the author or the moderator through PushShift [^2] https://pushshift.io.
 The statistics of the collected comments and submissions are shown in Appendix A.1.
-To balance the number of comments and submissions across different communities, for each community, we keep a maximum of 140K comments and 30K submissions.
-Therefore, there are at most 170K texts, which makes up the text corpus $D$, including comments and submissions, for each community.
+We do not differentiate between comments and submissions, and treat each comment and submission as a separate document.
+In subsequent sections of this paper, we use “comment” to refer to both “comment” and “submission”.
+4 Instructional Data Generation
+For each domain, we curate a collection of (1) instructional datasets CommInst for community-specific instruction tuning, which is used to finetune LLMs to views of different communities, and (2) survey datasets CommSurvey for evaluating the alignment of the finetuned LLMs’ views to the communities using survey questions.
+The pipeline is depicted in Figure 3.
 4.1 Topic Modeling
 For a domain, denote the combined corpus from all $n$ communities by $\mathbb{D}=D_{1}\cup D_{2}\cup....\cup D_{n}$.
-We use BERTopic Grootendorst (2022) on $\mathbb{D}$ to identify topics $g$.
+We use BERTopic Grootendorst (2022) on $\mathbb{D}$ to identify topics $T$.
 More details about BERTopic can be found in Appendix B.1.
-After topic modeling, each text $d$ is assigned a topic $g(d)$.
-The number of topics detected in politics and fitness are 425 and 628 respectively.
-The top 10 most frequent topics in politics and fitness are shown in Appendix B.1 respectively.
-Politics topics include abortion, climate, guns, and racism.
-Fitness topics include body measures, keto diet, and weight lifting.
-For each community $C_{i}$, we split the text corpus $D_{i}$ into smaller chunks $D_{i}=\{\hat{D}^{i}_{1},\hat{D}^{i}_{2},....\}$, where each chunk $\hat{D}$ consists of texts belonging to a single topic $g(\hat{D})$.
-3.3 Open-ended Instruction-Answer Pair Generation
-For a specific domain, starting with an empty demonstration pool, we randomly sample without replacement a text chunk $D_{i}$ for each topic $t_{k}\in T$ from each community.
-We include the sampled text chunks on topic $t_{k}$ from the $n$ communities into a prompt and then use the prompt to query the advanced LLM to generate open-ended instructions and answers on the topic.
+After topic modeling, each text $d$ is assigned a topic $t(d)$.
+For each community $C_{i}$, we split the text corpus $D_{i}$ into smaller chunks $D_{i}=\{\hat{D}^{i}_{1},\hat{D}^{i}_{2},....\}$, where each chunk $\hat{D}^{i}$ consists of 50 randomly sampled texts belonging to a single topic $t(\hat{D}^{i})$.
+A chunk $\hat{D}^{i}$ on topic $t(\hat{D}^{i})$ is considered an occurrence of the topic.
+To investigate different views across the communities, we only keep the topics that are discussed by at least $n-1$ communities.
+To balance the occurrences of different topics, for each topic in a community, we keep a maximum of 5 chunks.
+As a result, for politics and diet, there are 41 and 148 topics respectively.
+We present the keywords of 10 topics that appear 5 times in all communities in Appendix B.1.
+Politics topics include COVID vaccines, abortion, climate, guns.
+diet topics include body measures, keto diet, and eating disorder recovery.
+4.2 Open-ended Instructions Generation
+For a specific domain with $n$ communities, we initialize $n$ empty demonstration pools $\{\textsc{CommInst}_{i}\}_{i=1}^{n}$.
+For a topic $t_{k}\in T$ that is discussed by $m$ communities [^3] As we require a topic to appear at least $n-1$ communities, $m=n$ or $m=n-1$., from each community $C_{i}$, we randomly sample without replacement a text chunk $\hat{D}^{i}$ on $t_{k}$.
+We include the $m$ sampled text chunks $\{\hat{D}^{1},\hat{D}^{2},...,\hat{D}^{m}\}$ on topic $t_{k}$ into a prompt and then use the prompt to query the advanced LLM (GPT-4o) to generate open-ended instructions and responses on the topic.
 The prompting template is shown in Appendix B.2.
-We query claude-3-opus-20240229 using the Anthropic API.
 In the prompt, we first specify the topic by the topic keywords, and then the comments in the sampled chunks from different communities.
-Next we ask the model to generate instructions that can be answered based on the texts, and that the instructions should elicit different responses from the $n$ communities.
-Each generated instruction is paired with different answers from different communities (Figure 2(a)).
-It is then added to the pool as $n$ demonstrations.
-Note that some topics may be discussed more in certain communities than others, so some instructions may be paired with answers from less than $n$ communities.
-We iteratively repeat this process until there are no text chunks left for topic $t_{k}$ from any community, until we shift to the next topic $t_{k+1}$.
-After the iterative generation process, the instruction-answer pool will contain demonstrations covering various topics, which will use to finetune foundational LLMs.
-Please refer to Appendix A.2 for more details about CommInst.
-For fair evaluating the finetuned LLMs, we split CommInst-Pol in two ways:
-  • Random split. CommInst-Pol is randomly split into training and eval sets: CommInst-Pol-train and CommInst-Pol-eval, in a ratio of 85:15.
-  • Topical split.
-We randomly select 10 topics in politics and collect demonstrations related to the topics into CommInst-Pol-eval-topic.
-The remaining demonstrations are collected as CommInst-Pol-train-topic.
-In this way, CommInst-Pol-train-topic and CommInst-Pol-eval-topic has no topic overlap.
-We repeat the same process for CommInst-Fit.
-4 CommSurvey: A Collection of Community-Specific Survey Questions
+Next, we instruct the model to generate an instruction that can be answered based on the texts, and that the instruction should elicit different responses from the communities.
+In addition, we instruct the model to not rely on its pre-knowledge about the community, but solely focus on the given texts.
+The generated instruction $X$ is paired with different responses $\{Y_{i}\}_{i=1}^{m}$ from different communities, as shown in Figure 2(a).
+Each demonstration $\{X,Y_{i}\}$ is then added to the corresponding pool for the $i$-th community.
+We iteratively repeat this process until there are fewer than $n-1$ communities that have chunks left on topic $t_{k}$, then we shift to the next topic $t_{k+1}$.
+After the iterative generation process, each demonstration pool $\textsc{CommInst}_{i}$ is used to finetune a foundational LLM for community $C_{i}$.
+Please refer to Appendix B.3 for more details about CommInst.
+4.3 Multi-choice Surveys Generation
 To measure the alignment of the finetuned foundational LLM to the community, we administer a survey of multi-choice questions to the model and evaluate the agreement between the model’s responses and the community members’ responses.
 However, manually designing surveys and collecting human responses from online communities is a non-trivial process, which is costly and time-consuming.
-Instead, we curate another dataset called CommSurvey for evaluating the alignment of the finetuned LLM’s views to the community using survey questions.
-CommSurvey consists of automatically generated multi-choice survey questions and answers, where each question is paired with answers from different communities (Figure 2(b)).
-CommSurvey-Pol and CommSurvey-Fit are the two subsets for politics and fitness respectively, consisting of 4,457 and 2,930 multi-choice questions respectively, covering a wide range of topics.
-politics r/Anarchism
-r/Anarcho_
-Capitalism
-r/Conservative r/Libertarian r/Republican r/democrats
-0.75 0.60 0.75 0.45 0.75 0.65
-fitness r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
-0.75 0.70 0.60 0.50 0.65
-Table 1: Agreement between human annotators’ survey responses and the advanced LLM (Claude-3) generated survey answers in different domains and subreddits, measured by accuracy.
-4.1 Multi-choice Survey Generation.
-We initialize an empty multi-choice survey pool.
-Then, following the open-ended instruction-answer pairs generation in §$\S$3.3, we iteratively generate multi-choice questions and answers (Figure 2(b)) using the same advanced LLM (Claude-3) and add them to the pool.
-We also explored other advanced LLMs, such as GPT-4 or Gemini, but none of them was able to generate high-quality questions or answers as Claude-3.
-The prompting template is shown in Figure 8.
+Instead, we curate another collection of datasets called CommSurvey for evaluating the alignment of the finetuned LLMs’ views to the communities.
+We initialize $n$ empty multi-choice question pools $\{\textsc{CommSurvey}_{i}\}_{i=1}^{n}$.
+Then, following the open-ended instruction-response pairs generation process in §$\S$4.2, we iteratively generate multi-choice questions and answers (Figure 2(b)) using the same advanced LLM (GPT-4o) and add them to the pools [^4] In practice we use the advanced LLM to generate open-ended instructions and multi-choice questions at the same time in a single query.
+For clarity of presentation, we articulate them separately..
+The prompting template is shown in Appendix B.2.
+Each question pool $\textsc{CommSurvey}_{i}$ is used to evaluate a finetuned foundational LLM on community $C_{i}$.
+Please refer to Appendix B.3 for more details about CommSurvey.
 We assume that the answers generated by the advanced LLM are the “semi-ground truths”, and that they faithfully represent the views of the corresponding communities.
-Note that there is a correspondence between instructions in CommInst and survey questions CommSurvey, as they are generated using nearly identical prompts.
-Following the correspondence, we split CommSurvey-Pol and CommSurvey-Fit following the same two ways for CommInst.
-Due to the correspondence, there is no information leakage between CommInst-Pol-train and CommSurvey-Pol-eval, or between CommInst-Pol-train-topic and CommSurvey-Pol-eval-topic.
-4.2 Validating Generated Survey Answers
-We verify the faithfulness of the “semi-ground truths” for the communities via human annotation.
-For each domain, we randomly sample 20 questions from the survey (see Fig.
-9 and Fig.
-10 in the Appendix) and ask two expert annotators to answer them from the perspective of the subreddit.
-We also tried directly posting the surveys to the subreddits and having community members answer them.
-However, our posts were immediately removed by moderators and our accounts were banned from certain subreddits.
-This further indicates the difficulty of directly surveying online communities.
-[…]
+As an empirical evaluation, we compute the pairwise agreement between different communities using their responses to questions in CommSurvey.
+The results are presented in Figure 5 in Appendix B.4.
+We observe significant polarization between the left-leaning and right-leaning communities, and strong agreement between the two right-leaning communities (r/Conservative and r/AskThe_Donald).
+However, in the diet domain where the communities are more practical, there is less polarization.
+For more rigorous evaluation, we verify the “semi-ground truths” by human annotation, as detailed in Appendix C.
 5 Experiments
-We demonstrate the effectiveness of the pipeline on the Reddit politics and fitness domains.
-We align LLMs to Reddit communities by finetuning them on relevant demonstrations from CommInst.
-After finetuning, we administer surveys from CommSurvey and evaluate the finetuned LLMs by comparing their responses to the “semi-ground truths”, in order to measure alignment.
+We align foundational LLMs to Reddit communities by finetuning them on relevant demonstrations from CommInst.
+After finetuning, we evaluate the models using surveys from CommSurvey, comparing their responses to “semi-ground truths” to assess alignment.
 5.1 Experimental Setup
-We experiment with two foundational LLMs – GPT-3.5 and Llama-3.
+Data Generation.
+We use OpenAI’s GPT-4o Batch API to generate the datasets.
+For each query, three open-ended instructions and two multi-choice questions are created, as shown in Appendix B.2.
+The API costs approximately \$2 for the politics domain and \$5 for the diet domain.
+Data Splitting.
+Among all the queries used to prompt the advanced LLM to generate data, we randomly select 85% of them and use the instructions generated from these queries as the training instructions denoted as CommInst-train.
+For the rest 15% queries, we use the corresponding multi-choice questions as test questions, denoted as CommSurvey-test.
+In addition to randomly splitting the queries, we perform a topic-wise split, and make sure that the 85% training instructions, CommInst-train-topic, and 15% test questions, CommSurvey-test-topic, do not cover the same topics.
 Finetuning.
-We finetune the foundational LLMs on CommInst-Pol-train and CommInst-Fit-train.
-We use the OpenAI finetuning API to finetune gpt-3.5-turbo-0125 for 1 epoch, as we find that the loss converges within 1 epoch.
-The batch size is automatically determined by the API, and the finetuning takes ~40mins.
-For Llama-3, the implementation is based on LlamaFactory Zheng et al.
-(2024), and we finetune Llama-3-8B-Instruct with batch size 16 for 2 epochs on 4 H100 GPUs.
-The finetuning takes ~10mins.
+We focus on two strong foundational LLMs – Llama-3.1-8B-Instruct Dubey et al.
+(2024) and GPT-3.5-Turbo Ouyang et al.
+(2022).
+For each community $C_{i}$, we finetune the LLM on CommInst-traini (§$\S$5.3) and CommInst-train-topici (§$\S$5.4).
+The input and the output are the instruction and response verbatim.
+Llama-3.1 is finetuned with LlamaFactory Zheng et al.
+(2024), using both full and LoRA finetuning Hu et al.
+(2022), with batch size 16 for 3 epochs.
+We report the results of the model (full or LoRA) that achieves better loss on the validation set (5% of the training data).
+Full finetuning takes around 3 minutes on 4 NVIDIA H100 GPUs, and LoRA finetuning takes around 30 seconds on 1 GPU.
+For GPT-3.5, we use the OpenAI API, which completes finetuning in around 10 minutes for 1 epoch, where the batch size is automatically determined.
 Measuring Alignment.
-We administer survey questions in CommSurvey-Pol-veal and CommSurvey-Fit-veal to the finetuned foundational LLMs.
-In the prompt, we provide the question and the four options verbatim, and then add the instruction “Select only one answer by stating either A, B, C, or D.
-Do not provide any additional explanation or rationale for your choice.”, for the sake of easier matching of the model’s response to the option.
-For GPT-3.5, the temperature is set to 0.9 and the model provides 5 generations for each question.
-For Llama-3, temperature is set to 0.9, $top\_p$ is set to 0.9, and the model provides 10 generations for each question.
-The majority vote from the different generations for each question is taken.
-We compare the model responses to the “semi-ground truths” and report the accuracy as a measure of alignment.
+We administer survey questions from CommSurvey-test (§$\S$5.3) and CommSurvey-test-topic (§$\S$5.4) to the finetuned foundational LLMs.
+Each prompt includes the question and options verbatim, followed by the instruction, “Select only one answer by stating either A, B, C, or D.
+Do not provide any additional explanation or rationale for your choice,” to facilitate easier matching of the model’s responses to the options.
+We set the temperature to 0.8, generating 20 responses per question, and take the majority vote.
+Accuracy is calculated by comparing the model’s answers to the “semi-ground truths”.
+Subreddit Llama-3.1-8B GPT-3.5-Turbo
+Without Steering With Steering Without Steering With Steering
+Context CrossInst Steering Steering+CrossInst Context CrossInst Steering Steering+CrossInst
+r/Liberal 8.3 54.2 55.8 58.3 41.7 62.5 45.8 62.5
+r/NeutralPol 3.8 50.0 55.0 63.3 55.0 55.0 40.0 50.0
+r/Anar_Cap 54.1 76.7 70.0 66.7 50.0 66.7 66.7 73.3
+r/Conservative 18.4 63.3 60.5 56.7 50.0 53.3 53.3 53.3
+r/AT_Donald 23.7 56.7 66.7 70.0 30.0 56.6 50.0 63.3
+avg. politics 21.7 60.2 61.6 63.0 45.3 58.8 51.2 60.5
+r/keto 26.0 75.0 67.0 72.0 56.0 68.0 66.0 66.0
+r/WLAdvice 27.7 60.6 65.2 61.7 59.6 58.5 66.1 64.8
+r/EDAnony 29.1 72.1 65.1 69.8 61.6 72.1 62.7 66.3
+avg. diet 27.6 69.2 65.8 67.8 59.1 66.2 64.9 65.7
+Table 1: Evaluation results of Llama-3.1-8B and GPT-3.5-Turbo on CommSurvey-test.
+The community-aligned LLMs (CrossInst and Steering+CrossInst) are finetuned on CommInst-train, so there is potential topic overlap between training and evaluation.
+For each model family, the results are divided into two groups, one without and one with steering.
+The best results in each group are highlighted in bold.
 5.2 Baselines
-Vanilla LLMs.
-We administer survey questions to the vanilla unfinetuned GPT-3.5 and Llama-3 and evaluate their answers.
-Vanilla LLMs+Steering.
-When prompting the vanilla language models to answer survey questions, we steer them to mimic the community, by specifying in the prompt that “Select only one answer that best aligns with the opinions of members from subreddit r/[subreddit]”.
-Steering the vanilla LLMs can nudge them to respond for the community.
-However, it is important to note that steering only applies to predefined communities that are developed via a predefined tag, such as subreddits, where we can easily specify community names in the prompt to steer the LLMs.
-Although in this paper we focus on communities associated with online forums, Community-Cross-Instruct generalizes to other online communities, including organically formed communities in the retweet network Chu et al.
+LLM+Context.
+For each community $C_{i}$, we provide the vanilla (unfinetuned) LLM with the context about the community by appending 300 most relevant comments from it.
+For example, when prompting the LLM to answer a question in CommSurvey-testi, we retrieve the most relevant 300 comments to the question, from the text chunks that are used to generate CommSurvey-traini, by calculating the embedding similarity with sentence-transformers.
+The prompt is augmented with the instruction “According to the following statements, learn the mindset and select only one most relevant answer by stating either A, B, C, or D.
+Do not provide any additional explanation or rationale for your choice.
+[comments]”.
+This baseline is inspired by the idea of in-context learning, where the LLM learns the community’s mindset within the context provided in the prompt.
+Providing the LLM with context becomes less efficient when the model is deployed to answer a large number of survey questions, as each question requires processing an extremely long input sequence due to the added comments.
+This significantly increases computational costs, memory usage, and processing time, making it inefficient for large-scale or real-time applications.
+In contrast, our framework, which finetunes the model on community-specific data, is a one-time effort.
+Once the model is finetuned, it can efficiently answer survey questions without requiring additional long contextual inputs, making it far more scalable and resource-friendly in practical use cases.
+LLM+Steering.
+When prompting the vanilla LLM to answer survey questions, we steer it to mimic the community, by specifying in the prompt that “Select only one answer that best aligns with the opinions of members from subreddit r/[subreddit]”.
+Steering the vanilla LLMs can nudge them to respond to the community.
+For fair comparison to this baseline, we also steer the finetuned LLM via Community-Cross-Instruct, where the LLM is aligned to the community both in the finetuning and steering process.
+It is worth noting that steering only applies to predefined communities that are developed via a manually specified tag, such as subreddits, where we can easily reference community names in the prompt.
+While this paper focuses on forum-based communities, Community-Cross-Instruct generalizes to other online communities, including organically formed communities in the retweet network Chu et al.
 (2024) or the news co-sharing network Mosleh et al.
 (2021); He et al.
-(2024b), as long as their texts are readily available.
-It is non-trivial to summarize the organically-formed communities concisely in language, but we can still align LLMs to them via Community-Cross-Instruct.
+(2024b), as long as their relevant text data is readily available.
+Although it is not always feasible to concisely summarize organically formed communities in text, Community-Cross-Instruct allows for LLM alignment to such communities without requiring explicit textual descriptions.
+Subreddit Llama-3.1-8B GPT-3.5-Turbo
+Without Steering With Steering Without Steering With Steering
+Context CrossInst Steering Steering+CrossInst Context CrossInst Steering Steering+CrossInst
+r/Liberal 11.1 50.0 58.3 61.1 41.7 61.1 63.9 65.1
+r/NeutralPol 20.8 60.0 54.2 65.0 52.5 65.0 67.5 67.5
+r/Anar_Cap 23.9 56.5 52.2 56.5 54.3 58.7 58.7 71.7
+r/Conservative 20.0 54.3 52.5 65.2 50.0 67.4 69.6 65.2
+r/AT_Donald 25.0 56.5 57.5 56.5 52.2 60.9 61.2 63.0
+avg. politics 20.2 55.5 54.9 60.9 50.1 62.6 64.2 66.5
+r/keto 29.7 64.4 65.2 63.6 59.3 67.8 61.0 63.5
+r/WLAdvice 26.9 63.5 61.5 62.5 58.7 62.5 63.5 62.5
+r/EDAnony 25.4 55.9 50.0 54.9 57.8 60.8 58.8 61.8
+avg. diet 27.3 61.3 58.9 60.3 58.6 63.7 61.1 62.6
+Table 2: Evaluation results of Llama-3.1-8B and GPT-3.5-Turbo on CommSurvey-test-topic.
+The community-aligned LLMs (CrossInst and Steering+CrossInst) are finetuned on CommInst-train-topic, so there is no topic overlap between training and evaluation.
+For each model family, the results are divided into two groups, one without steering and one with steering.
+The best results in each group are highlighted in bold.
 5.3 Main Results
-We compare survey answers generated by the finetuned LLMs to the “semi-ground truth” and report the accuracy in Figure 4 as a measure of alignment with the corresponding community.
-In most cases, steering LLMs towards specific communities creates better alignment by relying on the models’ pre-existing knowledge about the communities.
-However, such pre-knowledge may be incomplete, inaccurate, or outdated.
-Therefore, by finetuning LLMs on instructions distilled from the text corpora of communities, Community-Cross-Instruct outperforms steering and achieves the best results, demonstrating its effectiveness in aligning LLMs to different communities.
-(a) politics
-(b) fitness.
-Figure 4: Alignment of finetuned LLMs to the corresponding communities measured by accuracy, by comparing the LLMs’ survey responses to the “semi-ground truths”.
-The base models show worse alignment with r/Conservative and r/Republican than other Reddit communities.
-This observation conforms to findings by previous works Santurkar et al.
-(2023); He et al.
-(2024a); Chen et al.
-(2024b) that LLMs tend to display a liberal bias and poor alignment with conservatives.
-After explicitly steering them towards specific communities, their alignment with the two subreddits significantly increases, but it is still lower than that with other communities.
-Finetuning the LLMs with Community-Cross-Instruct further improves the alignment, although it is still lower than for other communities.
-This observation highlights the risks of deploying LLMs in political scenarios due to the entrenched political biases.
-For the r/Anarchism subreddit, GPT-3.5-CrossInst achieves similar alignment as the steering baseline.
-We suspect that this is because views about anarchism are captured by the pre-existing knowledge of GPT-3.5.
-Therefore, infusing such knowledge again by Community-Cross-Instruct barely leads to any gain.
-For the r/Fitness and r/xxfitness communities, steering both LLMs worsens the alignment compared to the vanilla model.
-r/Fitness is a general fitness subreddit and r/xxfitness is a fitness community for women, both of which may have a broader and more diverse range of opinions compared to other subreddits.
-Steering the models to align with a singular “community opinion” might oversimplify the nuanced perspectives within these fitness communities, leading to decreased alignment accuracy.
-However, Community-Cross-Instruct effectively distills knowledge from the community corpora during finetuning.
-On r/GettingShredded, Community-Cross-Instruct is outperformed by the steering baseline for GPT-3.5, but not Llama-3.
-r/GettingShredded has the least amount of comments and posts, producing much fewer demonstrations for finetuning (Appendix 3).
-The insufficient data might be less effective for finetuning a larger LLM like GPT-3.5111We assume that GPT-3.5 has more than 8B parameters., leading to inferior performance.
+We compare the finetuned LLMs’ generated answers to multi-choice questions to the “semi-ground truths” and report the accuracy in Figure 1 as a measure of alignment with the corresponding community.
+The LLMs are finetuned on CommInst-train and evaluated on CommInst-test.
+Politics.
+In the politics domain, CrossInst consistently outperforms Context across both LLMs, demonstrating the strength of finetuning on community-specific instructional data.
+By aligning LLMs to explicit community instructions, CrossInst enables the models to better capture the nuances of political discourse, where ideological divides and subtle differences in language are critical.
+This makes the model more adept at accurately reflecting community values without being overwhelmed by noisy or redundant information, which is a common issue with Context.
+We observe that the a large portion of generated answers from Llama-3.1 using Context are texts irrelevant to the survey questions, indicating that it struggles in dealing with a lengthy prompt with 300 examples.
+An important aspect to consider is that LLMs come with pre-existing knowledge about various communities, which is learned during pretraining on large-scale internet data.
+During Steering, the model attempts to retrieve and apply this pre-existing knowledge to align its responses with the given community.
+However, this knowledge may not always be fully accurate or consistent with the actual values of the community being evaluated.
+For example, Steering might prompt the model to draw on generalizations or stereotypes that are present in its pretraining data, which might not reflect the current or specific views of the subreddit in question.
+This inconsistency can explain why Steering alone sometimes leads to suboptimal performance.
+When combining Steering with CrossInst, a potential conflict arises between the knowledge gained from finetuning and the pre-existing community knowledge the model tries to retrieve through steering.
+The model is essentially being pulled in two directions: one based on the explicit, finetuned instructions from CrossInst and the other based on its internalized, sometimes outdated, pretraining knowledge.
+This conflict can result in Steering+CrossInst underperforming compared to Steering alone, as observed in certain subreddits, such as r/Anarcho_Capitalism on Llama-3.1.
+In these cases, the inconsistency between the steering prompts and the finetuned knowledge creates confusion for the model, leading to lower performance.
+Diet.
+In the diet domain, we observe a similar pattern where CrossInst significantly outperforms the Context baseline.
+Communities like r/keto and r/EDAnonymous, which are focused on specific health and lifestyle goals, benefit greatly from instructional finetuning.
+These communities are characterized by practical and focused discussions, and CrossInst allows the model to adapt to the specific language and norms within these subreddits, ensuring a more accurate reflection of their values and objectives.
+Interestingly, in r/WeightLossAdvice, Steering alone performs the best across the four methods within each LLM.
+One possible reason for this is that in certain communities, particularly those that are well-represented in pretraining data, the model’s pre-existing knowledge might be more aligned with the actual community values.
+In such communities, where strong ideological markers may already be embedded in the model’s pretraining data, steering allows the model to retrieve this relevant information effectively, leading to stronger performance.
+In such cases, Steering helps amplify the model’s pre-learned alignment with the community, which can be sufficient for capturing the community’s voice without the need for additional finetuning.
 5.4 Out-of-Topic Generalizability
@@
 In the real world, the surveys may contain questions covering topics that do not appear in the finetuning data CommInst.
-In this study, we finetune Llama-3 on CommInst-Pol-train-topic, and evaluate it on CommSurvey-Pol-eval-topic, to make sure that there is no overlap in topic coverage.
-The results in Fig.
-5 show that Community-Cross-Inst consistently outperforms both baselines, demonstrating its superiority in aligning with views of different communities, and generalize the learned views to unseen topics.
-Figure 5: Cross-topic alignment of finetuned LLMs to the corresponding communities. The LLMs are finetuned on demonstrations and evaluated on survey questions on different topics.
+In this study, we finetune the LLMs on CommInst-train-topic, and evaluate them on CommSurvey-test-topic, to make sure that there is no overlap in topic coverage.
+The results are shown in Table 2.
+CrossInst models continue to outperform the Context baselines, demonstrating their strong generalization capability to new topics, suggesting that CrossInst maintains its robustness even when exposed to entirely new topics.
+We observe that Steering alone occasionally performs the best (e.g., r/AskThe_Donald on both LLMs, and r/WeightLossAdvice on GPT-3.5).
+We argue that this might be because ideologically-polarized or practical communities are consistent in their mindsets over time, so the LLM’s pre-knowledge about them would well predict their ideologies in the future.
 6 Related Works
-Self-Improved Instruction Tuning.
+6.1 Self-Improved Instruction Tuning.
 A survey of instruction-tuning approaches Zhang et al.
@@
 (2024c), and bootstrapping from an existing instruction set to generate new ones Wang et al.
-(2022).
+(2022); Chen et al.
+(2024b).
 Additionally, a human-in-the-loop framework builds upon self-generalization by iterating between human and machine-generated instructions Guo et al.
@@
 This work significantly contributes to the field by offering a scalable and fully autonomous solution to instruction tuning, paving the way for more adaptive and intelligent models.
-Community Alignment.
-Aligning with the dialects and behaviors of online communities is crucial for accurately modeling subpopulations Dorn et al.
+6.2 Aligning LLMs to Subgroups
+Existing work has aligned LLMs to different human subgroups to discover their mindset and opinions Dorn et al.
 (2023, 2024).
@@
 (2023a).
-Using self-instruct with questions from the American Trans Panel as the seed instruction set, LLMs can be ideologically manipulated to reflect particular political opinions Chen et al.
-(2024b).
+Argyle et al.
+(2023b) find that exposing GPT-3 to thousands of socio-demographic backstories leads the model to obtain a complex understandings of sociological concepts.
 To learn about partisan communities, Jiang et al.
-(2022) repeatedly prompt GPT 2 with tweets authored by prominent community members to generate synthetic responses to election survey questions.
-Argyle et al.
-(2023b) find that exposing GPT 3 to thousands of socio-demographic backstories leads the model to obtain a complex understandings of sociological concepts.
-‘Media diet models’ train BERT using MLM on news articles, evaluating the alignment of model predictions with a survey of community members who follow the media Chu et al.
-(2023).
-Feldman et al.
-(2022) expose GPT 3 to Yelp reviews to expand insights past sentiment information.
-Prompting language models with (survey question, individual ID, survey year) triplets from the Americans in the General Social Surveys has helped social scientists impute missing survey data and predict opinions Kim and Lee (2023).
-These techniques rely on manually created datasets at least as a seed dataset to align the models with the desired communities, limiting SRBMs to high-resource environments.
-Surveys.
-Social scientists use surveys to systematically collect data from populations to characterize their beliefs, attitudes, opinions, and behaviors.
-While designing the survey, researchers consider the order, wording, and format of questions (open-ended vs multiple-choice) to increase clarity and ensure the validity and reliability of the survey data.
-Survey design also aims to minimize biases, such as the non-response bias Hill et al.
-(1997), fatigue due to overly long surveys Choi and Pak (2005), social desirability bias Gordon (1987) when asking survey participants questions about sensitive topics (due to the desire to conform to perceived social expectations), or those involving social stigmas Goel and Salganik (2010).
-These issues, as well as the variety of existing biases, call for new approaches to survey design.
+(2022) propose CommunityLM by finetuning GPT-2 on tweets authored by prominent community members, and prompt the finetuned model to generate opinions about various political entities.
+He et al.
+(2024b) extends CommunityLM to organically-formed online communities with more fine-grained ideologies.
+However, these finetuned GPT-2 models can only be used for text continuation tasks and cannot answer survey questions.
+In our paper, we finetune the foundational instruction-tuned LLMs to serve as their digital twins.
+The resulting LLMs retain the instruction-following capabilities and are able to complete various tasks as specified in the instructions, including survey completion.
+6.3 Evaluating LLMs’ Subgroup Alignment
+Social scientists use surveys to systematically collect data from populations to characterize their beliefs, attitudes, opinions, and behaviors Hill et al.
+(1997); Choi and Pak (2005); Gordon (1987); Goel and Salganik (2010).
+LLM developers and practitioners focus on measuring LLM’s alignment with different human subgroups Santurkar et al.
+(2023); Durmus et al.
+(2023); He et al.
+(2024a) using use real-world survey responses as ground truth for comparing values.
+However, they often ignore under-represented groups, as it is difficult to administer surveys to people from those groups.
+To administer surveys, researchers sample individuals from a target population and ask them to manually respond to survey questions.
+This can be costly and slow, which limits the utility and timeliness of surveys.
+Our method addresses this challenge by using social media data that is easily accessible, even for those hard-to-reach subgroups.
+By automatically creating multi-choice survey questions and answers CommSurvey for these groups, our method enables a more comprehensive and cost-effective evaluation of LLMs’ alignment with diverse groups.
+7 Discussions
+We aim to align LLMs to online communities through instruction-tuning so that the aligned LLMs can be flexibly surveyed to probe the communities’ views on different topics.
+To prepare data for instruction tuning, we propose Community-Cross-Instruct, and the automatic creation of the open-ended instruction-response pairs (CommInst) is the core contribution.
+However, in order to demonstrate that the finetuned LLMs are indeed aligned to communities and can be treated as their digital twins, we propose to automatically generate survey questions and answers (CommSurvey) to efficiently evaluate LLM alignment.
+This is a secondary contribution of this paper, and we do not claim that CommSurvey is the best way to measure alignment.
+Therefore, our main goal is to propose a method to prepare LLMs as a proxy for surveying different online communities, rather than automatically creating such surveys.
 8 Conclusion
@@
 Through finetuning LLMs based on community-specific instructions and answers, Community-Cross-Instruct enables the generation of accurate responses that align with the beliefs and perspectives of different online communities.
-This innovative framework has demonstrated its effectiveness in representing various communities, including political and fitness forums on Reddit.
-Community-Cross-Instruct opens up new possibilities for researchers to gain insights into the diverse views and opinions present in online communities, paving the way for deeper understanding and analysis of digital societies.
+This innovative framework has demonstrated its effectiveness in representing various communities, including political and diet forums on Reddit.
+Community-Cross-Instruct opens up new possibilities for researchers to gain insights into online communities’ diverse views and opinions, paving the way for deeper understanding and analysis of digital societies.
 Moving forward, we will apply the framework to model organic online communities, such as those in the retweet network.
+In addition, we are interested in aligning LLMs to communities through reinforcement learning from human feedback (RLHF).
 Limitations
+Ignoring Thread Structure in Reddit Data.
+We treat each Reddit comment and submission as independent data points, without taking into account the thread structure in which they are embedded.
+Reddit discussions are inherently hierarchical, with comments often building upon or responding to previous comments and submissions.
+By ignoring this structure, we may lose important contextual information that could help capture the nuances of community interactions and opinions.
+Future work could explore leveraging the thread structure to better model the flow of conversations, capturing how community members engage with each other’s ideas and how opinions evolve within discussions.
 Identical LLM for Generating both Datasets.
-We use Claude-3 for generating both CommInst for finetuning LLMs and CommSurvey for evaluating LLM alignment, as we couldn’t find a different advanced LLM that was able to generate high-quality data as Claude-3.
+We use GPT-4o for generating both CommInst for finetuning LLMs and CommSurvey for evaluating LLM alignment, as we couldn’t find a different advanced LLM that was able to generate high-quality data as GPT-4o.
 This reliance on a single LLM introduces a limitation, as it may lead to a bias where the evaluation dataset (CommSurvey) is inherently aligned with the model used for instruction generation (CommInst), potentially overestimating the alignment accuracy of the finetuned models.
@@
 Additionally, we will include human-in-the-loop validation to ensure the generated datasets maintain high quality and representativeness, mitigating any biases introduced by the single LLM dependency.
-Furthermore, we will explore cross-validation techniques and third-party evaluations to benchmark the performance of our finetuned models, ensuring robustness and generalizability of our results beyond the influence of a single LLM.
+Furthermore, we will explore cross-validation techniques and third-party evaluations to benchmark the performance of our finetuned models, ensuring the robustness and generalizability of our results beyond the influence of a single LLM.
 Group Approximation.
-To approximate group-level behavior, community members represented in the minority are inherently excluded.
-Further, measuring group alignment using a single-answer multiple choice questionnaire does not account for bimodal distribution.
+To approximate group-level behavior, community members represented in the minority might be inherently excluded.
+Further, measuring group alignment using a single-answer multiple choice questionnaire does not account for a more complex distribution of the various opinions within the community.
 We hope to build on this work by experimenting with survey designs that account for more diversity in communities.
@@
 In this work, we work primarily with one prompt for instruction generation.
-We would be interested to see how model’s generated instructions shift with different prompting schemas.
+We would be interested to see how the model’s generated instructions shift with different prompting schema.
 Ethics Statement
+Finetuning LLMs towards Bias.
+Aligning LLMs with specific communities may result in models that appear more biased, as they are finetuned to the distinctive views and perspectives of those communities.
+However, this process is done solely for the purpose of accurately reflecting the values and attitudes prevalent within each community, rather than intentionally making the models more biased.
+Our goal is not to reinforce or amplify harmful biases, but rather to provide a computational tool that can represent the views of a given community for research purposes.
+To mitigate potential misuse, these models should be used in controlled environments and in contexts where understanding specific community perspectives is essential for social, cultural, or political research.
+User Privacy and Consent.
 Users might not be informed of how their data (reflected in their posts) are used to produce digital twins that mirror their voices and the purpose of the survey constructed based on their data.
 Furthermore, automatically surveying communities can reveal unconsented insights of certain individuals or groups of people online.
-To address these ethical concerns, we will implement several measures.
-Firstly, we will ensure that all data used for creating digital twins is anonymized, stripping any personally identifiable information to protect user privacy.
-Secondly, we will seek to aggregate data in a manner that focuses on community-level insights rather than individual-level analysis, thereby reducing the risk of unconsented personal exposure.
+To address these ethical concerns, we implement several measures.
+Firstly, we ensure that all data used for creating digital twins is anonymized, stripping any personally identifiable information to protect user privacy.
+Secondly, we seek to aggregate data in a manner that focuses on community-level insights rather than individual-level analysis, thereby reducing the risk of unconsented personal exposure.
 Thirdly, we will obtain explicit consent from users where possible, clearly communicating the purpose and scope of our research.
 Lastly, we will adhere to ethical guidelines and institutional review board (IRB) requirements to ensure that our methods respect the privacy and rights of the individuals and communities involved.
-Appendix A Data Details
+Acknowledgements
+We are grateful to Patrick Gerard, Ashwin Rao, and DJ Berry for helping review responses to surveys.
+This project was funded in part by the Defense Advanced Research Projects Agency (DARPA) under Agreement Nos.
+HR00112290021 and HR001121C0168.
+Appendix A Reddit Data
 A.1 Subreddits
@@
 Subreddit Description
-r/Anarchism
-r/Anarchism is a forum with 268K members for discussing anarchism, assuming
-participants are anarchists, allies, or learners.
+r/Liberal
+r/Liberal is a subreddit dedicated to discussions and news from a liberal political
+perspective, focusing on progressive policies, social justice, and left-leaning viewpoints.
+r/NeutralPolitics
+r/NeutralPolitics is a community dedicated to evenhanded, empirical discussion of
+political issues. It is a space to discuss policy and the tone of political debate.
 r/Anarcho_Capitalism
-r/Anarcho_Capitalism is a subreddit with 197K members dedicated to discussing
+r/Anarcho_Capitalism is a subreddit dedicated to discussing
 free-market capitalist anarchism and related topics, advocating for a society where
@@
 r/Conservative
-r/Conservative, with 1.1M members, offers the largest space on Reddit for fiscal
-and social conservatives to explore and discuss political and cultural issues from
-a distinctly conservative perspective.
-r/Libertarian
-r/Libertarian is a forum with 506K members focused on libertarianism and the
-principles of a free society, explicitly rejecting anti-libertarian ideologies
-like communism.
-r/Republican
-r/Republican is a partisan subreddit with 191K members specifically designed
-as a space for Republicans to discuss issues among themselves.
-r/democrats
-r/democrats is a subreddit with 438K members dedicated to supporting the
-Democratic Party, offering daily updates, policy analysis, and opportunities to
-help elect Democrats at all levels.
-r/Fitness
-r/Fitness is the largest community dedicated to achieving physical fitness
-goals with 12M members
-r/loseit
-r/loseit is a supportive community with 4M members for individuals of all
-sizes to discuss healthy and sustainable weight loss methods
+r/Conservative offers the largest space on Reddit for fiscal and social
+conservatives to explore and discuss political and cultural issues from a distinctly
+conservative perspective.
+r/AskThe_Donald
+r/AskThe_Donald is a subreddit that serves as a hub for supporters of former President
+Donald Trump, fostering discussions around conservative politics, pro-Trump views,
+and right-wing ideologies.
 r/keto
-r/keto is a community of 3.6M members for sharing experiences and advice
-about the low-carb Ketogenic diet, which supports a range of health
-conditions from diabetes to epilepsy
-r/xxfitness
-r/xxfitness is a community of 2.2M members for female and gender
-non-binary/gender non-conforming Redditors to discuss fitness
-r/GettingShredded
-r/GettingShredded is a community of 162K members for anyone aiming to lose
-body fat, improve their physique, and boost confidence through healthy fat loss
-and strength training discussions.
-Table 2: Reddit forums used in this study. Descriptions of (top) politics and (bottom) fitness subreddits.
-A.2 Statistics of CommInst
-Table 3 describes the number of demonstrations in each subreddit.
-The variance is due to their different topic coverage.
-Figure 6 presents the distribution of instruction lengths and answer lengths in CommInst-Pol and CommInst-Fit.
-Politics
-r/Anarch. r/Anarcho. r/Conserv. r/Democr. r/Libert. r/Republic.
-Comments 182,192 589,354 2,391,745 142,274 1,950,089 140,572
-Submissions 31,567 41,264 356,842 90,001 40,967 48,808
-Demonstrations 2,430 3,027 2,943 3,159 2,928 2,574
-Fitness
-r/Fitness r/loseit r/keto r/xxfitness r/GettingShredded
-Comments 907,839 967,645 603,920 420,082 78,053
-Submissions 2,430 3,027 2,943 3,159 2,928
-Demonstrations 3,438 2,793 1,716 2,883 1,215
-Table 3: Dataset size measured by number of comments and submissions for each community in political and fitness communities, and the number of learned open-ended instruction-answer pairs (demonstrations).
-(a) Instructions in CommInst-Pol
-(b) Instructions in CommInst-Fit
-(c) Answers in CommInst-Pol
-(d) Instructions in CommInst-Fit
-Figure 6: Distribution of instruction and answer lengths in CommInst-Pol and CommInst-Fit.
-Appendix B Implementation Details
+r/keto is a community for sharing experiences and advice about the low-carb Ketogenic
+diet, which supports a range of health conditions from diabetes to epilepsy
+r/WeightLossAdvice
+r/WeightLossAdvice is a subreddit where users share tips, strategies, and support for healthy
+and sustainable weight loss, with a focus on practical advice and personal experiences.
+r/EDAnonymous
+r/EDAnonymous is a subreddit that provides a supportive, anonymous space for individuals
+struggling with eating disorders to share their experiences, seek advice, and offer
+encouragement on the path to recovery.
+Table 3: Reddit forums used in this study.
+r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
+# of comments 31,233 35,725 670,686 2,243,842 142,543
+# of submissions 5,243 2,072 22,551 153,813 15,645
+r/keto r/WLAdvice r/EDAnonymous
+# of comments 603,466 225,635 347,477
+# of submissions 49,571 42,840 99,925
+Table 4: Number of comments and submissions in each subreddit.
+Appendix B Instructional Data Generation
 B.1 Topic Modeling with BERTopic
@@
 The embedding dimensions are reduced with UMAP, with n_neighbors 15 and n_components 5.
-Then the embeddings are clustered using HDBSCAN, with min_cluster_size 50.
+Then the embeddings are clustered using HDBSCAN, with min_cluster_size 40.
 After fitting the model, we remove texts that are not assigned any topics.
-Table 4 and Table 5 present the top-10 most frequent topics in politics and fitness.
-idx Topic Keywords Frequency
-1 covid, vaccinated, masks, vaccines, healthcare, deaths, flu, pandemic, coronavirus, cdc 48,787
-2 abortion, fetus, abortions, roe, body, choice, unborn, rights, parenthood, cells 11,256
-3 climate, change, solar, emissions, pipeline, fuels, coal, power, supply, grid 9,035
-4 guns, firearms, shootings, firearm, amendment, laws, armed, rifles, militia, nra 8,833
-5 police, officer, officers, gun, force, law, crime, video, training, situation 7,934
-6 kamala, woman, pelosi, vp, hillary, president, herself, biden, party, with 7,879
-7 racist, racism, whites, racists, racial, minorities, democrats, privilege, whiteness, discrimination 7,785
-8 libertarian, libertarians, libertarianism, party, government, left, ideology, vote, political, conservative 6,666
-9 anarchism, anarchy, anarchist, anarchists, anarcho, hierarchies, communism, ideology, government, system 6,311
-10 biden, bernie, obama, candidate, voters, debate, election, democrats, hillary, sanders 5,768
+Table 5 and Table 6 present the top 10 most frequent topics in politics and diet.
+idx Topic Keywords
+1 vaccine, covid, vaccinated, vaccines, flu, unvaccinated, vaccination, mrna, pandemic, data
+2 climate, prices, solar, change, co2, emissions, coal, cars, fuels, earth
+3 ballots, fraud, election, voter, ballot, evidence, voters, machines, audits, rigged
+4 abortion, abortions, fetus, roe, murder, unborn, conception, choice, womb, cells
+5 ukraine, russia, putin, nato, war, ukrainians, crimea, invade, conflict, eu
+6 twitter, facebook, musk, google, social, platforms, companies, users, censorship, ban
+7 gun, guns, firearms, shootings, firearm, rifles, laws, amendment, ammo, armed
+8 impeachment, mueller, fbi, impeach, documents, collusion, comey, congress, whistleblower, crimes
+9 gender, trans, transgender, lgbt, dysphoria, children, sexuality, identity, pronouns, genders
+10 capitol, blm, antifa, riots, riot, insurrection, rioters, protesters, terrorism, january
 Table 5: Top-10 most frequent topics in politics. Each topic is represented by the topic-10 keywords.
-idx Topic Keywords Frequency
-1 lbs, bmi, down, cw, gained, progress, healthy, tips, weeks, started 18,296
-2 keto, ketosis, carbs, diet, 20g, protein, macros, started, calories, ketogenic 13,078
-3 leggings, clothes, wear, pants, shorts, bras, shirts, clothing, dress, lululemon 7,063
-4 motivation, journey, progress, yourself, change, motivated, success, start, proud, back 5,142
-5 water, scale, weighing, retention, trend, normal, plateau, fluctuates, gain, weekly 4,880
-6 sleep, morning, hours, work, gym, motivation, schedule, waking, exercise, routine 4,486
-7 she, people, overweight, weight, obese, comment, healthy, skinny, talk, shaming 4,192
-8 bf, percentage, fat, bodyfat, body, lean, bulk, estimates, lbs, pic 3,930
-9 reps, sets, program, max, amrap, lifts, ranges, bench, 5x10, hypertrophy 3,788
-10 shoulder, elbow, injury, arm, knees, physio, exercises, bicep, curls, heal 3,775
-Table 5: Top-10 most frequent topics in fitness. Each topic is represented by the topic-10 keywords.
+idx Topic Keywords
+1 keto, protein, ketosis, started, macros, back, weeks, carbs, diet, calories
+2 skinny, thin, girls, like, overweight, body, underweight, weight, hate, myself
+3 thank, therapy, relapse, recover, help, life, support, yourself, proud, treatment
+4 scale, week, water, lose, weighing, daily, trend, fluctuates, months, plateau
+5 potassium, sodium, magnesium, electrolytes, mg, supplement, ketoade, citrate, 5000mg, chloride
+6 alcohol, drinking, beer, drink, drinks, liquor, alcoholic, gin, booze, drinker
+7 protein, macros, fat, min, kcal, 20g, carbs, bf, lean, need
+8 binging, restricting, binges, eating, hunger, control, bingeing, cycle, guilt, feeling
+9 ed, eds, recovery, recover, therapist, treatment, me, help, coping, talk
+10 coffee, tea, chocolate, brew, creamer, starbucks, latte, unsweetened, flavors, sweetener
+Table 6: Top-10 most frequent topics in diet. Each topic is represented by the topic-10 keywords.
 B.2 Prompting Template for CommInst
-The prompting template for generating the open-ended instruction-answer pairs in CommInst is shown in Figure 7.
-Figure 7: Template prompt to generate CommInst.
-B.3 Prompting Template for Generating CommSurvey
-The prompting template for generating the multiple choice instruction-answer pairs in CommSurvey is shown in Figure 8.
-Figure 8: Template prompt to generate CommSurvey.
-Appendix C Survey Sample
-Samples of the surveys provided to human annotators for generating “semi-ground truth data” are provided in Figures 9 and 10 (political and fitness domains, respectively).
-Figure 9: Sample of the multiple choice survey generated for human evaluation in the politics domain.
-Figure 10: Sample of the multiple choice survey generated for human evaluation in the fitness domain.
+The prompting template for generating the open-ended instruction-response pairs in CommInst and multi-choice question-answer pairs in CommSurvey is shown in Figure 4.
+Figure 4: Prompting template to generate CommInst and CommSurvey in the politics domain.
+B.3 Statistics of CommInst and CommSurvey
+Table 7 describes the number of open-ended instruction-response pairs in CommInst and multi-choice question-answer pairs in CommSurvey. The variance is due to their different topic coverage.
+r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
+# of instructions 234 234 300 303 303
+# of survey questions 156 156 200 202 202
+r/keto r/WLAdvice r/EDAnonymous
+# of instructions 1,032 885 870
+# of survey questions 688 590 580
+Table 7: Number of generated open-ended instruction-response pairs in CommInsti and multi-choice question-answer pairs in CommSurveyi, for each community.
+B.4 Community Agreement Analysis
+For a pair of communities $C_{i}$ and $C_{j}$, we identify their common survey questions in CommSurveyi and CommSurveyj, and compute their agreement on the questions using Cohen’s Kappa.
+The agreement matrices for politics and diet are shown in Figure 5.
+In the politics domain, we observe significant polarization between communities.
+For example, the agreement between r/Liberal and both r/Conservative (0.01) and r/AskThe_Donald (0.03) is extremely low, indicating the deep ideological divide between these left-leaning and right-leaning communities.
+This low agreement reflects how vastly different their responses are to the same political survey questions, aligning with the broader patterns of political polarization seen across online platforms.
+On the other hand, r/Conservative and r/AskThe_Donald exhibit much higher agreement (0.59), highlighting the ideological overlap between these conservative-leaning communities, likely reflecting their shared values and viewpoints on key political issues.
+Meanwhile, r/NeutralPolitics shows moderate agreement with both r/Anarcho_Capitalism (0.24) and r/Liberal (0.28), suggesting that as a centrist or neutral forum, it shares certain perspectives with both libertarian and liberal ideologies.
+Additionally, r/Anarcho_Capitalism shows higher agreement with conservative-leaning subreddits, such as r/Conservative (0.36) and r/AskThe_Donald (0.44), reflecting shared economic or libertarian principles, despite divergences on other political issues.
+In the diet domain, the relationships between the communities are more practical and less polarized compared to politics.
+Communities like r/keto and r/WeightLossAdvice exhibit moderate agreement (0.39), likely reflecting shared goals related to weight loss and health, even though they may emphasize different strategies.
+Similarly, r/EDAnonymous and r/keto (0.38) show moderate overlap, suggesting that while their focus areas are different—one being centered on eating disorders and the other on ketogenic diets—they share some common ground in their approaches to health and diet-related topics.
+The strongest alignment within the diet domain is between r/WeightLossAdvice and r/EDAnonymous (0.41), which likely stems from overlapping concerns about diet and body image, which are central to both communities’ discussions.
+(a) politics
+(b) diet
+Figure 5: Pairwise agreement between different communities, measured by Cohen’s Kappa.
+Appendix C Human Annotation
+We verify the faithfulness of CommSurvey as the “semi-ground truths” for the communities via human annotation.
+We also tried directly posting the survey questions to the subreddits and having community members answer them.
+However, our posts were immediately removed by moderators and our accounts were banned from certain subreddits.
+This further indicates the difficulty of directly surveying online communities.
+C.1 Annotator Recruiting
+The four annotators for the politics domain closely follow political news and are knowledgeable about the platforms of the parties represented in our sample.
+The three annotators for the diet domain are active members of diet and diet and have substantial knowledge of diet-related topics and trends.
+To ensure the annotators were capable of accurately answering the survey questions for different communities, we implemented a thorough selection and training process.
+Annotators were chosen based on their demonstrated expertise and familiarity with the relevant domains.
+Additionally, we conducted preliminary tests where annotators answered a subset of questions, and their responses were evaluated for consistency and accuracy.
+C.2 Annotation Process
+We randomly sampled 20 questions from CommSurvey-Politics (Figure 6) and 10 questions from CommSurvey-Diet (Figure 7) for the annotators to answer.
+When answering each question, the annotators were instructed to search the relevant posts from the subreddit and learn their views from the discussions, instead of solely relying on their pre-assumption about the community.
+Each annotator filled out the question for a different subset of the communities, as shown in Table 8.
+If multiple annotators annotated for a community, they had another round of discussions to resolve their disagreement.
+As a result, for community, the annotators delivered one set of annotations of the survey questions.
+Figure 6: Ten multiple choice survey questions (out of the twenty) in CommSurvey for human evaluation in the politics domain.
+Figure 7: Ten multiple choice survey questions CommSurvey for human evaluation in the diet domain.
+C.3 Annotation Evaluation
+r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
+Annotator 1 x x
+Annotator 2 x x x x
+Annotator 3 x x
+Annotator 4 x x
+r/keto r/WLAdvice r/EDAnonymous
+Annotator 1 x x x
+Annotator 2 x x
+Annotator 3 x x
+Table 8: Communities that each annotator annotated for.
+For each subreddit, we compare the LLM-generated “semi-ground truths” to the label from human annotators, and present the accuracy in Table 9.
+In both domains, “semi-ground truths” achieve strong agreement with human annotations for most subreddits, which gives confidence that the advanced LLM-generated survey answers can be used as “semi-ground truth” to evaluate finetuned foundational LLMs.
+r/Lib r/NeutralPol r/Anarcho_Cap r/Conserv r/ATDonald
+Human-LLM Acc 0.75 NA 0.65 0.55 0.55
+r/keto r/WLAdvice r/EDAnonymous
+Human-LLM Acc 0.7 0.6 0.5
+Table 9: Agreement between human annotators’ survey responses and the advanced LLM (GPT-4o) generated survey answers, measured by accuracy.